# SoundMorpher: Perceptually-Uniform Sound Morphing with Diffusion Model

### v1

@@ 4.3 Controllable Sound Morphing with Discrete $\alpha$ Series
Static morphing. To achieve controllable static morphing, we control the target SPDP point $p$, which represents how the desired output perceptually intermediate between $x^{(0)}$ and $x^{(1)}$. We find the corresponding $\alpha$ value by the binary search with the target $p$ and further obtain a morphed result ${x^{(\alpha)}}$. Pseudocodes for static morphing are in Algorithm 3.
@@
Cyclostationary morphing. To produce $N$ perceptually uniform hybrid sounds between $x^{(0)}$ and $x^{(1)}$, we first obtain $N$ uniform interpolated SPDP points $\{p_{i}\}^{N}_{i=1}$. Then we find corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ and further obtain $N$ morphed results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ as in Algorithm 4.
@@
Dynamic morphing. Dynamic morphing performs sound morphing over time, but one challenge is that if the morphing path fails to ensure perceptual intermediateness and content correspondence, the resulting sounds may exhibit perceptual discontinuities or unnatural intermediate stages. As in Algorithm 5, we obtain $N$ interpolated target SPDP points $\{p_{i}\}^{N}_{i=1}$ with $\Delta p$. The corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ are determined by binary search with the target SPDP points. Each morphed result $x^{(\alpha_{i})}$ contributes a segment of duration $\frac{T}{N}$, producing an audio segment $\tilde{x}^{(\alpha_{i})}$ according to index $i$. The final audio signal is obtained by concatenating these morphed segments, resulting in
$[x_{0},x_{1},...,x_{T}]=\text{concat}(\tilde{x}^{(0)},\tilde{x}^{(\alpha_{1})},...,\tilde{x}^{(1)})$ (8)
@@
Table 1: Timbral morphing for musical instruments compared to baseline on different instruments.
@@
5 Experiment
In this section, we showcase three applications of SoundMorpher in real-world scenarios: Timbral morphing for musical instruments, Environmental sound morphing, and Music morphing.
5.1 Evaluation Metric
@@
5.2 Timbral morphing for musical instruments
Sound morphing can allow timbral morphing between the sound of two known musical instruments, creating sounds from unknown parts of the timbre space (McAdams, 2013; McAdams & Goodchild, 2017). Timbral morphing for musical instruments involves transitioning between timbres of two different musical instruments to create a new sound. This new sound could possess characteristics of both original timbres as well as new timbral qualities between them, which usually applied to creative arts. In this experiment, we perform timbral morphing for isolated musical instruments given two recordings of the same musical composition played by different music instruments.
Dataset. To ensure high quality of paired composition musical instrument on timbral study, we selected 22 paired musical instrument composition samples from demonstration pages of musical timbre transform projects, MusicMagus (Zhang et al., 2024) and Timbrer (Kemppinen P., 2020). The paired samples have durations varies from 5s to 10s, with 16.0kHz and 44.01kHz, which involve 5 groups of instrument pairs: 2 paired samples of piano-violin; 10 paired samples of piano-guitar; 1 paired sample of taiko-hihat; 1 paired sample of piano-organ, and 8 paired samples of harp-kalimaba.
@@ 5.2 Timbral morphing for musical instruments
Baseline. We compare our method with Sound Morphing Toolbox (SMT) (Caetano, 2019), which is a set of Matlab functions targeting on musical instrument morphing that implement a sound morphing algorithm for isolated musical instrument sounds. Since SMT does not offer guidance for selecting perceptually uniform morph factors, we uniformly interpolate 11 morph factors in $[0,1]$.
Results and analysis. We set $N=11$ for SoundMorpher with an initial prompt ‘a music composition by {instrument}’. The comparison of our method and the baseline on timbral morphing is in Table 1. Overall, SoundMorpher demonstrates superior morphing quality compared to STM across various metrics, including audio quality, intermediateness, smoothness, and resynthesis quality   [^2] Since we perform timbral morphing within the same music composition, MFCCsE may not a suitable metric under the same musical content. In contrast, we focus on evaluating smoothness and intermediateness., when applied to different types of musical instrument timbre morphing. Notably, STM fails in Taiko-Hihat timbral morphing due to significant high reconstruction perceptual error. In contrast, SoundMorpher maintains robustness across different types of musical instruments, making it a more flexible and efficient solution for timbral morphing applications on different types of musical instruments. Figure. 1 provides a visualization of normalized timbre space, illustrating morphing trajectories generated by SMT and SoundMorpher. The timbre space is defined by three important timbral features: Log-Attack Time, Spectral Centroid, and Spectral Flux (McAdams et al., 1995; McAdams, 2013). The SMT trajectory shows distinct steps, indicating that the transitions between each intermediate sound are relatively abrupt. The spacing between the blue points suggests that each step represents a significant change in timbre, which may result in a less smooth perceptual transition between two musical instruments. In contrast, the trajectory produced by SoundMopher demonstrates a smoother curve. The points are more closely spaced, indicating more gradual changes between each intermediate timbre. This suggests that SoundMopher achieves a more continuous and natural-sounding morphing process, with each step being a smaller, more refined adjustment compared to SMT. Figure 7 and Figure 8 in the appendix provides additional visualization for this experiment.
5.3 Environmental Sound Morphing
@@
Dataset. We use ESC50 (Piczak, 2015) which consists of 5-second recordings organized into 50 semantic classes which loosely arranged into 5 major categories. We randomly select 4 major categories of scenarios to verify our method, including (1) Dog-Cat (animals voices), (2) Laughing-Crying baby (human sounds), (3) Church bells-Clock alarm (urban noise-interior sound), (4) Door knock-clapping (interior sounds-human sounds). Each category of scenarios contains 25 randomly selected audio pairs, thereby, 100 randomly paired samples in total.
Table 2: Environmental sound morphing with different types of environmental sounds.
@@
[…]
5.4 Music Morphing
Dataset. In this experiment, we randomly selected 50 sample pairs from 20 musical samples available on AudioLDM2 (Liu et al., 2024) demonstration page. These 10-second music compositions that span different genres and feature both single or mixed musical instrument arrangements.
Figure 2: Visualization of dynamic morphed music with $N=15$, source music and target music.
Results and analysis. In this experiment, we select an initial text prompt as ‘a sound clip of music composition’ to perform dynamic morphing. Even though this experiment contains morph complex music compositions with different music genres and music instruments, Table 3 shows our method still superiors on perceptual smoothly transiting source music to the target music and ensures correspondence, intermediateness and smoothness. Figure 2 provides strong visual evidence that the dynamic morphing method effectively transitions from the source to the target music while maintaining perceptual smoothness, correspondence, and intermediate transformations. The spectrogram illustrates that the morphed music transitions gradually, maintaining smooth spectral changes over time, which suggests the method successfully morphs the source into the target music.
Table 3: Music morphing experimental results & ablation study for sound perceptual features, where N represents the number of components of PCA for reducing dimension of Mel-spectrogram.
[…]
Reduced Mel-Spec. (N=2) 0.187 10.31 58.57 0.793 0.056 $\pm$ 0.075 0.182
Reduced Mel-Spec. (N=3) 0.151 10.76 59.39 0.779 0.055 $\pm$ 0.079 0.151
Mel-Spec. 0.056 9.85 56.09 0.847 0.068 $\pm$ 0.045 0.178
MFCCs 0.053 10.11 57.38 0.987 0.071 $\pm$ 0.050 0.156
Spectral contras 0.066 10.54 58.44 0.863 0.061 $\pm$ 0.071 0.155
Table 4: Mean opinion score study on environmental sound morphing and music morphing task.
[…]
Environmental sound morphing 3.78 $\pm$ 0.31 3.67 $\pm$ 0.40 3.57 $\pm$ 0.47 3.67 $\pm$ 0.39
Music morphing 3.81 $\pm$ 0.39 3.55 $\pm$ 0.51 3.49 $\pm$ 0.48 3.62 $\pm$ 0.46
5.5 Discussion
Mean opinion score (MOS) study. We conducted a MOS study as a subjective evaluation for the correspondence, intermediateness, and smoothness of morphed results from SoundMorpher. The study involved 21 volunteers, and detailed methodology is in Appendix 9. As shown in Table 4, the results suggest that SoundMorpher is versatile, performing similarly well across both music and environmental sound morphing tasks, with no significant differences observed in the overall MOS. This consistency in scores indicates that SoundMorpher effectively handles the unique challenges posed by the distinct characteristics of music and environmental sounds, such as the continuous nature of music compared to the more segmented structure of environmental sounds. Despite the objective metric results showing clear differences between the two tasks, the human evaluation suggests that SoundMorpher remains robust across different sound types. One possible interpretation is that our objective metrics are more sensitive to variations in the measured aspects than participants.
Table 5: Comparison with MorphFader
@@ 5.5 Discussion
Model comparison. We compare SoundMorpher with a concurrent work, MorphFader (Kamath et al., 2024), based on the criteria outlined in Section 3.1. MorphFader relies on a pre-trained AudioLDM (Liu et al., 2023) and perform sound morphing by text instructions. We compare with 7 examples provided on the demonstration of MorphFader, that MorphFader uniformly samples $5$ morph factors in $[0,1]$, resulting in a morph path with $[0,0.25,0.5,0.75,1]$. In contrast, SoundMorpher finds $\alpha$ values according to constant $\Delta p$ with $5$ uniformly interpolate $p$ points by binary search. As in Table 5, SoundMorpher produces a smoother and perceptual intermediates morphing than MorphFader. See Appendix 10 for details.
Ablation study on sound perceptual features. We verified the perceptual feature in SPDP in music morphing task. We select alternative music information retrieval features (MIR) including MFCCs with 13 coefficients (Logan et al., 2000), and spectral contrast (Jiang et al., 2002). We use principal component analysis (PCA) to reduce the dimensionality of Mel-spectrogram to further capture variation of spectral content over time, which is referred to as reduced Mel-Spec. (Stevens et al., 1937; Casey et al., 2008; Jiang et al., 2002). Table 3 shows performance comparisons of SoundMorpher with different features, SoundMorpher with Mel-spectrogram achieves better morphing quality in terms of correspondence and smoothness variation with smaller FAD, FD and CDPAMstd. While Mel-spectrogram yields higher CDPAMmean, CDPAMT and MFCCsE compared to reduced Mel-Spec. and MFCCs, the differences in metric values are not significant. However, the overall morph quality with Mel-spectrogram is consistently better than other features. This suggests Mel-spectrogram, as a pseudo-3D representation, provide more perceptual and semantic information, which contributes to improve morph quality compared to higher-level features.
@@
Table 6: Experiment and ablation study results on music morphing.

### v2

@@ IV-D Controllable Sound Morphing with Discrete $\alpha$ Series
Static morphing. To achieve controllable static morphing, we control the target SPDP point $p$, which represents how the desired output perceptually intermediate between $x^{(0)}$ and $x^{(1)}$. We find the corresponding $\alpha$ value by the binary search with the target $p$ and further obtain a morphed result ${x^{(\alpha)}}$.
@@
Cyclostationary morphing. To produce $N$ perceptually uniform hybrid sounds between $x^{(0)}$ and $x^{(1)}$, we first obtain $N$ uniform interpolated SPDP points $\{p_{i}\}^{N}_{i=1}$. Then we find corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ and further obtain $N$ morphed results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$.
@@
Dynamic morphing. Dynamic morphing performs sound morphing over time, one challenge is that if the morphing path fails to ensure perceptual intermediateness and audio description correspondence, the resulting sounds may exhibit perceptual discontinuities or unnatural intermediate stages. SoundMorpher produces perceptually uniform sound morphing sequences, which can avoid the problem when performing dynamic morphing. We obtain $N$ interpolated target SPDP points $\{p_{i}\}^{N}_{i=1}$ with $\Delta p$. The corresponding morph factors $\{\alpha_{i}\}_{i=1}^{N}$ are determined by binary search with the target SPDP points. Each morphed latent variable $z^{(\alpha_{i})}$ contributes a segment of length $\frac{K}{N}$, where $K$ is the length of latent variables, producing a latent variable segment $\tilde{z}^{(\alpha_{i})}$ according to index $i$. The final morphed audio sample can be obtained by
$x_{morph}=\text{vocoder}(\text{decoder}(\text{concat}(\tilde{z}^{(\alpha_{1})},...,\tilde{z}^{(\alpha_{N})})))$ (10)
@@
TABLE I: Timbral morphing for musical instruments compared to the baseline on different instruments.
@@
V Experiment
In this section, we show three applications of SoundMorpher in real-world scenarios: Timbral morphing for musical instruments, Environmental sound morphing, and Music morphing.
V-A Evaluation Metric
We systematically verify SoundMorpher according to the criterions mentioned in Section III-A, with a series of adapted quantitative objective metrics as below
  • Correspondence. We design a metric that computes absolute error for the mid-point MFCCs proportion, namely MFCCsE, for descriptive correspondence. Let $N$ be an odd integer, We define a series of perceptually uniform morphing results $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ with source and target audio $x^{(0)}$ and $x^{(1)}$, where $i$ in the range of $1$ to $N$. The MFCCsE is computed by
    $(\displaystyle\text{MFCCs}_{\mathcal{E}}=\text{abs}(~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$11)
    $\frac{||m^{(\frac{N+1}{2})}-m^{(0)}||_{2}}{||m^{(\frac{N+1}{2})}-m^{(0)}||_{2}+||m^{(\frac{N+1}{2})}-m^{(1)}||_{2}}$ 0.5-0.5)- 0.5 )
    where $m^{(i)}$ represents the extracted MFCCs feature of the $i^{th}$ morphed results in the series $x^{(\alpha_{i})}$, $m^{(0)}$ and $m^{(1)}$ represents MFCCs feature of $x^{(0)}$ and $x^{(1)}$. This metric aims to evaluate spectrogram coherent of the midpoint result $x^{(\frac{N+1}{2})}$ between two end points $x^{(0)}$ and $x^{(1)}$, which MFCCs captures low-level of audio descriptive information [59, 60]. Ideally, we wish the midpoint morphed result contains half-and-half audio description elements on two end points. The larger $\text{MFCCs}_{error}$ indicates the audio description consistency is far away than the midpoint (i.e., 0.5). We extract MFCCs feature with 13 coefficients to compute MFCCsE.
    We use Fréchet audio distance (FAD) [61] and Fréchet inception distance (FID) [62] between morphed audios and sourced audios to verify semantic distribution similarity between morphed samples and sourced audios. Both metrics capture high-level of semantic audio descriptive information by extracting audio features from pretrained audio classification models. Given two sets of audio $X_{mrp}$ and $X_{src}$, where $X_{mrp}$ contains consecutive morphed samples as $X_{mrp}=\{x^{(\alpha_{i})}\}_{i=1}^{N}$, and $X_{src}$ contains sourced audio samples  [^1] In environmental sound morphing task, we set all sourced sounds in that class as $X_{src}$.. We calculate FAD and FID values based on audio features extracted by pretrained audio classification models between $X_{src}$ and $X_{mrp}$  [^2] The FAD and FID metrics are obtained based on https://github.com/haoheliu/audioldm_eval.
@@ V-A Evaluation Metric
  • Smoothness. We calculate the mean and standard deviation of CDPAM along with the morphing path to validate smoothness, as ${}_{mean}\pm\text{CDPAM}_{std}$, where $\text{CDPAM}_{mean}=\frac{1}{N-1}\sum^{N-1}_{i=1}\text{CDPAM}(x^{(\alpha_{i})},x^{(\alpha_{i+1})})$, and $\text{CDPAM}_{std}=\sqrt{\frac{1}{N-1}\sum^{N-1}_{i=1}(\text{CDPAM}(x^{(\alpha_{i})},x^{(\alpha_{i+1})})-\text{CDPAM}_{mean})^{2}}$. In timbre study, we define timbral distance as $\mathcal{L}_{2}^{timbre}$ by $\mathcal{L}_{2}^{timbre}=\frac{1}{N-1}\sum_{i=1}^{N-1}||q^{(\alpha_{i+1})}-q^{(\alpha_{i})}||_{2}$, where $q^{(\alpha_{i})}$ represents the corresponding timber point of $x^{(\alpha_{i})}$ in timbre space [66].
  • Reconstruction of perceptual correspondence. Lastly, we denote CDPAME that calculate CDPAM between $\{x^{(0)},x^{(1)}\}$ and $\{\hat{x}^{(0)},\hat{x}^{(1)}\}$, where $\hat{x}$ represents resynthesized end points when $\alpha=0$ and $\alpha=1$.
V-B Timbral Morphing for Musical Instruments
Sound morphing can allow timbral morphing between the sound of two known musical instruments, creating sounds from unknown parts of the timbre space [67, 68]. Timbral morphing for musical instruments involves transitioning between timbres of two different musical instruments to create a new sound. This new sound could possess characteristics of both original timbres as well as new timbral qualities between them, which usually applied to creative arts. In this experiment, we perform timbral morphing for isolated musical instruments given two recordings of the same musical composition played by different music instruments.
Dataset. To ensure high quality of paired compositions on timbral study, we selected 22 paired musical instrument composition samples from demonstration pages of musical timbre transform projects, MusicMagus [30] and Timbrer [69]. The paired samples have durations varies from 5s to 10s, with 16.0kHz and 44.01kHz, which involve 5 groups of instrument pairs: 2 paired samples of piano-violin; 10 paired samples of piano-guitar; 1 paired sample of taiko-hihat; 1 paired sample of piano-organ, and 8 paired samples of harp-kalimaba.
@@ V-B Timbral Morphing for Musical Instruments
Figure 3: Visualization of timbre morphing for musical instruments with $N=11$ by SoundMorpher compared to SMT.
Baseline. We compare with Sound Morphing Toolbox (SMT) [11], a set of Matlab functions targeting on musical instrument morphing. SMT implements a sound morphing algorithm for isolated musical instrument sounds. Since SMT performs sound morphing based on the linear assumption [16], we uniformly interpolate 11 morph factors in $[0,1]$.
Results and analysis. The comparison of our method and the baseline on timbral morphing is in Table I. Overall, SoundMorpher demonstrates superior morphing quality compared to STM across various metrics, including audio quality, intermediateness, smoothness, and resynthesis quality   [^3] Since we perform timbral morphing within the same music composition, MFCCsE may not a suitable metric under the same musical content. In contrast, we focus on evaluating smoothness and intermediateness., when applied to different types of musical instrument timbre morphing. Notably, STM fails in Taiko-Hihat timbral morphing due to significant high reconstruction perceptual error. In contrast, SoundMorpher maintains robustness across different types of musical instruments, making it a more flexible and efficient solution for timbral morphing applications on different types of musical instruments. It is not surprising that SMT has poor performance in this experiment, since the traditional signal processing sound morphing methods are limited to complex or inharmonic sounds [12]. Figure. 2 provides a visualization of normalized timbre space, illustrating morphing trajectories generated by SMT and SoundMorpher. The timbre space is defined by three important timbral features: Log-Attack Time, Spectral Centroid, and Spectral Flux [66, 67]. The SMT trajectory shows distinct steps, indicating that the transitions between each intermediate sound are relatively abrupt. The spacing between the blue points suggests that each step represents a significant change in timbre, which may result in a less smooth perceptual transition between two musical instruments. In contrast, the trajectory produced by SoundMopher demonstrates a smoother curve. The points are more closely spaced, indicating more gradual changes between each intermediate timbre. This suggests that SoundMopher achieves a more continuous and natural-sounding morphing process, with each step being a smaller, more refined adjustment compared to SMT. Fig. 3 provides spectrogram visualization of how SoundMorpher performs timbre morphing.
V-C Environmental Sound Morphing
@@
Dataset. In this experiment, we firstly use ESC50 [70] to study the robuteness of SoundMorpher accross different categories of environmental sounds. Then we make a direct comparison with a concurrent method, MorphFader [13], based on environmental sounds provided in  [^4] https://pkamath2.github.io/audio-morphing-with-text/webpage/index.html. ESC50 [70] which consists of 5-second recordings organized into 50 semantic classes which loosely arranged into 5 major categories. We randomly select 4 major categories of scenarios to verify our method, including (1) Dog-Cat (animals voices), (2) Laughing-Crying baby (human sounds), (3) Church bells-Clock alarm (urban noise-interior sound), (4) Door knock-clapping (interior sounds-human sounds). Each category of scenarios contains 25 randomly selected audio pairs, thereby, 100 randomly paired samples in total. To have a fair comparison, we also perform environmental sound morphing based on samples provided by MorphFader, which is sourced from AudioPairBank [71].
Baseline. Due to lack of open-sourced direct comparison, we compare SoundMorpher with MorphFader [13] based on the criteria outlined in Section III-A. Although MorphFader didn’t release their codes, our evaluation system still can make a fair comparison based on their demonstrated examples.
TABLE II: Environmental sound morphing with different types of environmental sounds.
@@
Figure 4: Visualization of spectrogram for morphed results compred with MorphFader. SoundMorpher appears to provide a more seamless and stable morphing process, in which transitions are smoother and the spectral content is more consistent across the morphing stages.
TABLE III: Comparison with MorphFader
@@
Figure 5: Visualization of environmental sound morphing with $N=5$, from top to bottom: (1) church bells $\leftrightarrow$ clock alarm (2) crying baby $\leftrightarrow$ laughing (3) cat $\leftrightarrow$ dog (4) clapping $\leftrightarrow$ wood door knocking
[…]
We then make a direct comparision with the concurrent method, MorphFader [13], based on their demonstrated morphed samples in Table III  [^5] Due to the demonstration page doesn’t provide original audio samples, we cannot compute FAD, FID and CDPAME for this comparison.. SoundMorpher produces a smoother and intermediate morph than MorphFader with improvement on CDPAMT and CDPAMmean. Fig. 4 provides a direct visual comparison of a morph sequence produced by MorphFader and our method. The yellow rectangles in Fig. 4 highlight frequency band intensity cross morphing results. The intensity of the frequency bands within the yellow rectangle changes more abruptly for MorphFader, in contrast, SoundMorpher are more stable and consistent across time. This suggests that SoundMorpher maintains better spectral consistency during the morphing process, with smoother transitions between different timbral characteristics. As red rectangles indicate, MorphFader introduces more abrupt changes at the end of the morphing sequence. There is a noticeable shift in the pattern, indicating less smoothness on transition. In contrast, SoundMorpher shows a more gradual and consistent transition within the red rectangles. The spectral patterns remain more stable and exhibit smoother transitions towards the end of the morphing sequence. Overall, SoundMorpher appears to provide a more seamless and stable morphing process. The transitions are smoother, and the spectral content is more consistent across the morphing stages.
V-D Music Morphing
Dataset. Motivated by [30], we randomly selected 50 sample pairs from 20 high-quality musical samples available on AudioLDM2 [42] demonstration page. These 10-second music compositions that span different genres and feature both single or mixed musical instrument arrangements.
Baseline. Due to lack of direct comparison on dynamic music morphing, we set the baseline as SoundMorpher without SPDP binary search, which achieves sound morphing by smoothly interpolate condition embeddings and latent states. We uniformly sample 15 morph factors in $\alpha\in[0,1]$ as the baseline for comparison.
Figure 6: Visualization of dynamic morphed music for SoundMorpher and baseline with $N=15$, source music and target music.
Results and analysis. Even though this experiment contains morph complex music compositions with different music genres and music instruments, Table IV shows our method superiors on perceptual smoothly transiting source music to the target music and ensures correspondence, intermediateness and smoothness compared to the baseline, which further illustrates the advantage of the proposed algorithm of SPDP with binary serach. Fig. 6 visually demonstrates the effectiveness of the proposed dynamic morphing method in transitioning smoothly from source to target music while ensuring perceptual consistency and intermediate transformations. Compared to the baseline results, which exhibit noticeable abrupt changes and spectral discontinuities, SoundMorpher produces gradual and smooth spectral transitions over time. This highlights that simply linearly interpolating semantic features, as done in the baseline method, fails to achieve truly smooth and perceptually consistent sound morphing.
@@ V-D Music Morphing
TABLE IV: Music morphing experimental results, baseline comparison & ablation study for sound perceptual features
[…]
Baseline (Mel-Spec.) 0.099 10.82 58.20 0.938 0.067 $\pm$ 0.065 0.158
SoundMorpher (Mel-Spec.) 0.047 9.85 56.25 0.847 0.068 $\pm$ 0.045 0.178
Reduced Mel-Spec. (n=2) 0.187 10.31 58.57 0.793 0.056 $\pm$ 0.075 0.182
Reduced Mel-Spec. (n=3) 0.151 10.76 59.39 0.779 0.055 $\pm$ 0.079 0.151
MFCCs 0.053 10.11 57.38 0.987 0.071 $\pm$ 0.050 0.156
Spectral Contrast 0.066 10.54 58.44 0.863 0.061 $\pm$ 0.071 0.155
$\star$ where $n$ represents the number of components of PCA for reducing dimension of Mel-spectrogram
V-E Discussion And Ablation Study
Ablation study on sound perceptual features. We verified the perceptual feature in SPDP in music morphing task. We select alternative music information retrieval features (MIR) including MFCCs with 13 coefficients [60], and Spectral contrastt [72]. We use principal component analysis (PCA) to reduce the dimensionality of Mel-spectrogram to further capture variation of spectral content over time, which is referred to as reduced Mel-Spec. [55, 73, 72]. Table IV shows performance comparisons of SoundMorpher with different features, SoundMorpher with Mel-spectrogram achieves better morphing quality in terms of correspondence and smoothness variation with smaller FAD, FID and CDPAMstd. While Mel-spectrogram yields higher CDPAMmean, CDPAMT and MFCCsE compared to reduced Mel-Spec. and MFCCs, the differences in metric values are not significant. However, the overall morph quality with Mel-spectrogram is consistently better than other features. This suggests Mel-spectrogram, as a pseudo-3D representation, provide more perceptual and semantic information, which contributes to improve morph quality compared to higher-level features.
TABLE V: Experiment and ablation study results on music morphing.